# PhoenixOS: Concurrent OS-level GPU Checkpoint and Restore with Validated Speculation

### v3

@@ Introduction
Fault tolerance for GPU processes (checkpoint-mostly).    Distributed computing applications such as training models are susceptible to GPU failures (DBLP:conf/icse/GaoSLZWLY23, ; DBLP:conf/nsdi/EisenmanMIMKNSA22, ; DBLP:journals/corr/abs-2202-07848, ). PhOS provides transparent fault tolerance through periodic GPU process checkpointing using our soft copy-on-write (CoW) protocol described in §4.2. The checkpoints are stored in fast fault-tolerant storage, such as with replications stored in remote memory (DBLP:conf/sosp/WangJZZFNW23, ). If a failure occurs, PhOS first stops all GPU processes, and then restores them from the most recent checkpoint. The checkpoint frequency is determined by holistically considering both the checkpoint overhead as well as the loss of computation due to restoring from a stale checkpoint, similar to prior work (DBLP:conf/eurosys/GuptaKKVGKRS24, ). The detailed method is left in the appendix (§A.1). For fault tolerance, using a slightly stale checkpoint (i.e., one that does not capture the absolute latest execution state) is acceptable, because the latest state can be recovered through recomputation.
Two things to note. First, fault tolerance is a checkpoint-mostly case, as the checkpoint frequency needs to be high to avoid losing computation results (DBLP:conf/sosp/WangJZZFNW23, ). In comparison, the restore only happens upon failures, which occur much less frequently, e.g., one per-hour (DBLP:conf/fast/MohanPC21, ; DBLP:journals/corr/abs-2205-01068, ). Second, we need to ensure the checkpoint from all the involved processes is consistent (DBLP:journals/corr/abs-2202-07848, ). Thus, we extended the quiescing phase described in §4.2 across all involved processes. After the quiesce, we can checkpoint each process with CoW separately. We currently follow Singularity (DBLP:journals/corr/abs-2202-07848, ) and use a user-provided hint for a correct global quiesce.
Table 2. Applications evaluated. indicates a multi-GPU setup. PPO is training-only, and our setup cannot run Llama3.3-70B training.
@@
Application Train Inference Train Inference Train Inference Train Inference Inference
@@
Fast GPU serverless function startups (restore-mostly).    In serverless computing, processes are created on demand for each request (DBLP:journals/corr/abs-1902-03383, ; ali-serverless-gpu, ). Since each process has an entry function to react to a request, we take a checkpoint right before the entry, so once a request comes, we can restore the process from the checkpoint using our concurrent restore protocol described in §6 to bypass the cold start cost (DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/eurosys/AoPV22, ) before the entry.
@@
Baselines.   We compare PhOS with NVIDIA’s official OS-level checkpoint and restore tool—cuda-checkpoint (nvidia-new-ckpt-driver-repo, ), which adopts a stop-the-world approach. Nevertheless, we found it is extremely slow, e.g., it cannot fully utilize the PCIe bandwidth to copy data, see Figure 11. Unfortunately, we don’t have its source code to analyze. To compare with the best performance of the stop-the-world C/R, we implemented Singularity (DBLP:journals/corr/abs-2202-07848, )—the state-of-the-art GPU C/R system, in our codebase as the baseline. We have carefully tuned the implementation, e.g., we leverage pinned memory to achieve maximum data copy performance. As shown in Figure 11, our implementation of Singularity has orders of magnitude smaller application stall time than cuda-checkpoint. Thus, we will omit a detailed comparison with cuda-checkpoint in later analysis. Finally, a recent work CRIUgpu (DBLP:journals/corr/abs-2502-16631, ) builds on cuda-checkpoint to support a unified stop-the-world checkpoint and restore. Its objective is for functionalities but not performance. Hence, we omit a comparison because it has the same performance as cuda-checkpoint.
Figure 12. The wasted GPU time using different checkpoint methods for fault tolerance on training workloads. The reported wasted time is normalized to the total time used for training ($T$), and the displayed time is normalized to the number of systems with the maximum wasted time for ease of interpretation. cuda-checkpoint does not support checkpointing distributed jobs.
Figure 13. The downtime of migrating processes between machines with different systems. Multi-GPU workloads are marked with .
Evaluated applications.   We focus on evaluating AI applications—the dominated GPU applications to show the effectiveness of PhOS. Table 2 summarizes the evaluated applications. We choose workloads of both training and inference, whose domains span vision tasks (ResNet), large foundational models (Llama2-13B and Llama3.3-70B), and reinforcement learning (PPO). These tasks may either use one or multiple GPUs. We follow the common setups for running these applications, and leave the detailed setup descriptions in the appendix (§A.3) due to space limitation.
@@
Fault tolerance: evaluated metrics and results.    To evaluate the effectiveness of PhOS in fault tolerance scenarios, we choose training applications. Following existing works (DBLP:conf/sosp/WangJZZFNW23, ; DBLP:conf/eurosys/GuptaKKVGKRS24, ), we compare different systems with the following metrics: checkpoint overhead and wasted GPU time during training. The checkpoint overhead measures the stall time caused by checkpointing, while the wasted GPU time measures the end-to-end training GPU time wasted when compared to an unrealistic, non-fault execution. For all systems, the checkpoint is stored in host memory to avoid slow storage (DBLP:conf/sosp/WangJZZFNW23, ).
@@
Figure 12 further compares the average wasted GPU hours caused by different checkpoint methods. Since the wasted time is related to the failure frequency and the checkpoint frequency, we set a failure ratio of one GPU failure per hour reported by various industry reports (DBLP:journals/corr/abs-2205-01068, ; DBLP:conf/fast/MohanPC21, ), and then calculate the optimal checkpoint frequency for each system according to the formula in §A.1. Note that different systems require distinct checkpoint frequencies: PhOS achieves optimal performance with 279 checkpoints per hour, whereas Singularity performs best at 67. PhOS can set a higher frequency thanks to the reduced checkpoint overhead. The results show that PhOS saves up to 22–86% GPU hours compared with other approaches: the higher checkpoint frequency of PhOS saves wasted GPU hours due to recomputation, while our concurrent checkpoint minimizes wasted GPU time caused by high checkpoint overhead with high checkpoint frequency. Interestingly, a well-implemented OS-level GPU C/R even outperforms a simple user-level checkpointing (torch.checkpoint) that uses PyTorch’s save.
Live migration: evaluated metric and results.    To evaluate the effectiveness of PhOS in live migration scenarios, we measure the downtime when migrating applications between nodes with different systems. Figure 13 shows the timeline of application performance during migration. We can see that PhOS incurs a minimal downtime thanks to the concurrent execution capabilities. Notably, PhOS only incurs 3.3 and 3.7 seconds of downtime for migrating Llama2-13B training and LLama2-70B inference, respectively. In comparison, Singularity incurs 10.2 and 12.35 seconds of downtime, which is dominated by copying data from the source GPU to the target. In Llama2-13B training, copying data through 100 Gbps RDMA takes at least 9.8 seconds.
Figure 14. The serverless function execution time under coldstart. The Exe. states for the function execution time. All the workloads are inference.
Figure 15. An analysis of the runtime validator overhead on (a) training and (b) inference workloads. (c) Analysis of the ratios of the kernels instrumented with the validator. -T is abbreviated for training while -I is for inference.
@@
Figure 16. A breakdown of the checkpoint overhead on (a) PhOS CoW, (b) without prioritized PCIe transfer optimization and (c) Singularity. Workload: Llama2-13B Training. Note that we omitted the concurrent data copy of PhOS.
@@
Effectiveness of the prioritized application PCIe transfer.    As shown in Figure 16, without our prioritized application PCIe transfer optimization described in §5 (b), applications still suffers stalls due to PCIe bandwidth competition. This is because the GPU kernels are waiting to load the training set from the CPU, which is blocked by the checkpoint data copy due to the limited number of GPU DMA engines on the GPUs.
Figure 17. A breakdown of the checkpoint overhead on (a) PhOS recopy, (b) without coordinated CPU GPU checkpoint and (c) Singularity. Workload: Llama2-70B Inference. Note that we omitted the concurrent data copy of PhOS.
Breakdown of the recopy protocol.    Figure 17 shows the breakdown of executing our recopy protocol on a Llama2-70B inference workload. We can see that the downtime of PhOS is dominated by the recopy time. Compared to CoW, the recopy downtime is longer because it not only ensures that the checkpoint is correct, but it is also fresh. Nevertheless, the downtime is still orders of magnitude smaller than a stop-the-world approach (Singularity) because the recopied data is much smaller (2.1 vs. 9.7 seconds).
@@
Figure 18. A breakdown of the restore on (a) PhOS concurrent restore, and (b) Singularity. Workload: Llama2-13B inference. Note that we omitted the concurrent data copy of PhOS. R stall states for restore stall time.
@@
The runtime overhead of tracing GPU read and write set.    Figure 15 (a) and (b) report the runtime overhead of the runtime validator on different applications. We observe a relatively small slowdown of 1–12% for various workloads. The observable overhead is small because: (1) the additional check only happens when the kernel accesses the global memory, which is less frequent for high computational efficiency; and (2) the instrumented kernels are only a small portion of the overall kernels in the workloads. Figure 15 (c) reports the number of instrumented kernels invoked during the concurrent checkpoint and restore: We can see that most of the kernels are not instrumented, e.g., in Llama2-13B inference, only 12% of the invoked kernels are instrumented.
@@
Figure 19. An analysis of the copy-on-write overhead with different on-device buffer sizes.
Figure 19 shows the impact of the on-device buffer size on the checkpoint overhead when checkpointing a Llama2-13B training workload. We choose this workload because it has the largest memory footprint such that little room is left for the on-device buffer (see Table 4). We report the increased stall time caused by the checkpoint, where the time is normalized to the stall time with our maximum configurable on-device buffer size (2 GB). We can see that checkpoint overhead generally decreases with the on-device buffer size, but the difference is small, e.g., even without on-device buffer, PhOS incurs up to 9.1 % more stalls. This is because not all copy-on-write operations require stalling the kernel, so the slow copy through PCIe with insufficient on-device buffer can be hidden.
@@
Figure 20. Impact of checkpoint timing on the performance of Llama2-13B training. The blank in App is the stalled caused by the CoW. The red indicates the data copy time. Note that we omitted the concurrent data copy of PhOS.
Figure 20 analyzes how the checkpoint timing affects the performance of PhOS on the CoW protocol. We omit analysis on the recopy protocol as it is similar. We control the timing with our SDK (see §A.2). We run LLama2-13B training workload and choose two times: (1) at the beginning of an iteration (before the forward pass), (2) at the end of the iteration (at the update phase). We measure the performance specifically on the 32th iteration to fully warm up the application. We can see that checkpoint at (1) is more efficient because few buffers are updated at the beginning: only activation buffers are updated (see also §8.5). Thus, if we can finish the checkpoint before reaching the update pass, we meet few stalls due to CoW. In (1), copy 2.3 GB of data via 32 GBps PCIe [^1] The measured bandwidth is 25 GBps, slightly below the hardware limit. takes 185 ms, while the iteration time is 6.3 s.
@@
Figure 21. A heatmap showing the read and write sets traced by our speculated validation method on a Llama2-13B training workload.
Figure 21 reports the GPU accesses traced by PhOS on Llama2-13B training. First, we can see that though PhOS traces accesses at buffer level, it is still fine-grained enough to support an efficient CoW and recopy, because GPU applications typically allocate and write buffers in a fine-grained way. Moreover, the access patterns differs across different phases of execution, so the checkpoint timing is important.
A study of the feasibility of speculation
Table 3. A study of the success rate of our speculation-based GPU read and write sets tracing beyond our evaluated applications.
GPU Apps #kernels/failed #Instances/failed
@@ Acknowledgment
We would like to thank the reviewers from SOSP’24, OSDI’24 and SOSP’25 for their insightful feedback. We sincerely thank Zhiyuan Dong, Dingyan Zhang and Xiating Xie for giving us feedback on the early draft of this paper. We would also like to thank Shaoxun Zeng for sharing details on how PhOS can be adapted to support CUDA graph. This work was supported in part by the National Key Research & Development Program of China (No. 2022YFB4500700), the Fundamental Research Funds for the Central Universities, the National Natural Science Foundation of China (No. 62202291, 62272291), as well as a research grant from Huawei. Corresponding author: Rong Chen (rongchen@sjtu.edu.cn).
@@ Appendix A Appendix
Table 4. Detailed setups of applications evaluated in §8. indicates a multi-GPU setup. PPO is training-only, and our setup cannot run Llama3.3-70B training.
@@
Application Train Inference Train Inference Train Inference Train Inference Inference
@@
# GPU used 1 1 1 N/A 8 1 8 1 N/A 8
Total GPU Memory (per GPU) 1.8 GB 1.7 GB 5.9 GB N/A 70.6 GB 8.9 GB 73.6 GB 55.4 GB N/A 70.8 GB
# GPU buffers (per GPU) 209 195 75 N/A 445 234 413 347 N/A 718
# Active GPU kernels 13 8 41 N/A 51 50 36 74 N/A 73
@@
Problem formulation.  Suppose we have $N$ GPUs involved in a distributed computing job, and each GPU fails $F$ times within an hour. Suppose each GPU failure is independent and identically distributed (i.i.d.), following a uniform distribution over the entire computation interval ($T$). The checkpoint overhead is $O$ with PhOS. If a GPU fails, PhOS will stop all the GPU processes, and restore from the latest checkpoint. The restore time is $R$ for all the GPUs.
The total GPU hour wasted due to failure is at a checkpoint frequency $f$ times per hour:
$\Bigl{(}N\,F\,T\Bigr{)}\;\times\;\Bigl{(}R\;+\;\tfrac{N}{2f}\Bigr{)}.$
@@
Put it all together: to total GPU hours wasted due to checkpoint overhead and failure recovery is:
$\text{Total GPU-hours wasted}\;=\;N\,F\,T\Bigl{(}R\;+\;\tfrac{N}{2\,f}\Bigr{)}\;+\;N\,O\,f\,T.$
Solving the optimal $f$.  For a given job, the distribution of failures ($F$), the checkpoint overhead ($O$), and the restore time ($R$) are static and can be profiled online (DBLP:conf/fast/MohanPC21, ), so we only need to solve the checkpoint frequency $f$ to minimize the total GPU hours wasted. By differentiating the total GPU hours wasted with respect to $f$, we got:
@@
Figure 22. An illustration of how applications can use PhOS SDK to control the checkpoint timing. Lines with + are code interacting with PhOS.
Applications can explicitly call our SDK to control the timing of checkpointing and the protocol used, see Figure 22. Thanks to our OS-level design, the code change is minimal: it only needs one line of code. In our example, the code tells PhOS that it needs to checkpoint the process at the beginning of an iteration of a training GPU process. Note that the checkpoint is asynchronously called so it will not block application execution unless the last checkpoint is not done.
@@
Table 4 shows the detailed buffer allocation information as well as the number of GPU kernels information of our evaluated applications. The detailed setups of training workloads are as follows: For ResNet50 training, we use the CIFAR-10 dataset and configure a batch size of 32 for training. For PPO training, we use the gym code (gym, ). For Stable Diffusion training, we use the SD v1-4 model and a batch size of 1,536 for each GPU. Finally, for Llama training, we use a distributed configuration of 8TP (Tensor Parallelism) and 1DP (Data Parallelism), the batch size is 4, restricted by the available GPU resources. All the training workloads use the AdamW optimizer.

### v4

@@ Introduction
Fault tolerance for GPU processes (checkpoint-mostly).    Distributed computing applications such as training models are susceptible to GPU failures (DBLP:conf/icse/GaoSLZWLY23, ; DBLP:conf/nsdi/EisenmanMIMKNSA22, ; DBLP:journals/corr/abs-2202-07848, ). PhOS provides transparent fault tolerance through periodic GPU process checkpointing using our soft copy-on-write (CoW) protocol described in §4.2. The checkpoints are stored in fast fault-tolerant storage, e.g., with replication to remote memory (DBLP:conf/sosp/WangJZZFNW23, ). If a failure occurs, PhOS first stops all GPU processes, and then restores them from the most recent checkpoint. The checkpoint frequency is determined by holistically considering both the checkpoint overhead as well as the loss of computation due to restoring from a stale checkpoint, similar to prior work (DBLP:conf/eurosys/GuptaKKVGKRS24, ). The detailed method is left in the appendix (§A.1). For fault tolerance, using a slightly stale checkpoint (i.e., one that does not capture the absolute latest execution state) is acceptable, because the latest state can be recovered through recomputation.
Two things to note. First, fault tolerance is a checkpoint-mostly case, as the checkpoint frequency needs to be high to avoid losing computation results (DBLP:conf/sosp/WangJZZFNW23, ). In comparison, the restore only happens upon failures, which occur much less frequently, e.g., one per-hour (DBLP:conf/fast/MohanPC21, ; DBLP:journals/corr/abs-2205-01068, ). Second, we need to ensure the checkpoint from all the involved processes is consistent (DBLP:journals/corr/abs-2202-07848, ). Thus, we extended the quiescing phase described in §4.2 across all involved processes. After the quiescent point, we can checkpoint each process with CoW separately. We currently follow Singularity (DBLP:journals/corr/abs-2202-07848, ) and use a user-provided hint for a correct global quiesce, e.g., before the forward pass of each training iteration.
@@
ResNET-152M PPO-336M Stable-Diffusion-1B Llama2-13B Llama3.3-70B
[…]
Application Training Inference Training Training Inference Training Inference Inference
@@
Fast GPU serverless function startups (restore-mostly).    In serverless computing, processes are created on demand for each request (DBLP:journals/corr/abs-1902-03383, ; ali-serverless-gpu, ). Since each process has an entry function to react to a request, we take a checkpoint right before the entry, so once a request comes, we can restore the process from the checkpoint using our concurrent restore protocol described in §6 to avoid the cold start costs (DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/eurosys/AoPV22, ) before entry.
@@
Baselines.    We compare PhOS with NVIDIA’s official OS-level checkpoint and restore tool—cuda-checkpoint (nvidia-new-ckpt-driver-repo, ), which adopts a stop-the-world approach. Nevertheless, we found it is extremely slow, e.g., it cannot fully utilize the PCIe bandwidth to copy data, see Figure 11. Unfortunately, we don’t have its source code to analyze. To compare with the best performance of the stop-the-world C/R, we implemented Singularity (DBLP:journals/corr/abs-2202-07848, )—the state-of-the-art GPU C/R system, in our codebase as the baseline. We have carefully tuned the implementation, e.g., we leverage pinned memory to achieve maximum data copy performance. As shown in Figure 11, our implementation of Singularity has orders of magnitude smaller application stall time than cuda-checkpoint. Thus, we will omit a detailed comparison with cuda-checkpoint in later analysis. Finally, a recent work CRIUgpu (DBLP:journals/corr/abs-2502-16631, ) builds on cuda-checkpoint to support a unified stop-the-world checkpoint and restore. Its objective is for functionalities, not performance. Hence, we omit a comparison because it has the same performance as cuda-checkpoint.
Figure 12. The comparison of process migration downtime between machines. () indicates a multi-GPU setup.
Evaluated applications.   We focus on evaluating AI applications—the dominant GPU applications to show the effectiveness of PhOS. Table 2 summarizes the evaluated applications. We choose workloads of both training and inference, whose domains span vision tasks (ResNET-152M), large foundational models (Llama2-13B and Llama3.3-70B), generation tasks (Stable-Diffusion-1B) and reinforcement learning (PPO-336M). These tasks may either use one or multiple GPUs. We follow common setups for running these applications, and leave the detailed setup descriptions in the appendix (§A.3).
@@
Fault tolerance: evaluated metrics and results.    To evaluate the effectiveness of PhOS in fault tolerance scenarios, we choose training applications. Following existing works (DBLP:conf/sosp/WangJZZFNW23, ; DBLP:conf/eurosys/GuptaKKVGKRS24, ), we compare different systems with the following metrics: checkpoint overhead and wasted GPU time during training. The checkpoint overhead measures the stall time caused by checkpointing, while the wasted GPU time measures the end-to-end training GPU time wasted when compared to a non-faulty execution. This includes both the ratio of checkpoint overhead within the overall execution as well as the recomputation time caused by recovering from a stale checkpoint. For all systems, the checkpoint is stored in host memory to avoid slow storage (DBLP:conf/sosp/WangJZZFNW23, ).
Figure 13. The wasted GPU time (proportional to the total training time $T$) using different checkpoint methods for fault tolerance for training workloads. To simplify comparison between applications, for an application, the bar is normalized to the number of systems with the maximum wasted time. cuda-checkpoint does not support checkpointing distributed jobs.
@@
Figure 13 further compares the average wasted GPU time caused by different checkpoint methods. Since the wasted time is related to the failure frequency and the checkpoint frequency, we set a failure ratio of one GPU failure per hour reported by various industry reports (DBLP:journals/corr/abs-2205-01068, ; DBLP:conf/fast/MohanPC21, ), and then calculate the optimal checkpoint frequency for each system according to the formula in §A.1. Note that different systems require distinct checkpoint frequencies: PhOS achieves optimal performance with 279 checkpoints per hour, whereas Singularity performs best at 67. PhOS can set a higher frequency thanks to the reduced checkpoint overhead. The results show that PhOS saves up to 22–86% GPU time compared with other approaches: the higher checkpoint frequency of PhOS saves wasted GPU time due to recomputation, while our concurrent checkpoint minimizes wasted GPU time caused by high checkpoint overhead with high checkpoint frequency. Interestingly, a well-implemented OS-level GPU C/R even outperforms a simple user-level checkpointing (torch.checkpoint) that uses PyTorch’s save.
Live migration: evaluated metric and results.    To evaluate the effectiveness of PhOS in live migration scenarios, we measure the downtime when migrating applications between nodes with different systems. Figure 12 shows the timeline of application performance during migration. We can see that PhOS incurs minimal downtime thanks to the concurrent execution capabilities. Notably, PhOS only incurs 3.3 and 3.7 seconds of downtime for migrating Llama2-13B training and Llama3.3-70B inference, respectively. In comparison, Singularity incurs 10.2 and 12.35 seconds of downtime, which is dominated by copying data from the source GPU to the target. In Llama2-13B training, copying data through 100 Gbps RDMA takes at least 9.8 seconds.
Figure 14. The breakdown of serverless function execution time in cold starts. Note that all workloads are inference.
@@
Figure 15. The breakdown of the checkpoint overhead on (a) PhOS CoW, (b) without prioritized PCIe transfer optimization, and (c) Singularity. Workload: Llama2-13B Training. Note that we omitted the concurrent data copy of PhOS.
@@
Effectiveness of the prioritized application PCIe transfer.    As shown in Figure 15, without our prioritized application PCIe transfer optimization described in §5, applications suffer from stalls due to PCIe bandwidth competition. This is because GPU kernels are waiting to load the training set from the CPU, which is blocked by the checkpoint data copy due to the limited number of GPU DMA engines.
Figure 16. The breakdown of the checkpoint overhead on (a) PhOS recopy, (b) without coordinated CPU and GPU checkpoint, and (c) Singularity. Workload: Llama3.3-70B Inference. Note that we omitted the concurrent data copy of PhOS.
Breakdown of the recopy protocol.    Figure 16 shows the breakdown of executing our recopy protocol on a Llama3.3-70B inference workload. We can see that the downtime of PhOS is dominated by the recopy time. Compared to CoW, the recopy downtime is longer because it not only ensures that the checkpoint is correct, but it is also fresh. Nevertheless, the downtime is still orders of magnitude smaller than a stop-the-world approach (Singularity) because the recopied data is much smaller (2.1 vs. 9.7 seconds).
@@
Figure 17. The breakdown of the restore on (a) PhOS concurrent restore, and (b) Singularity. Workload: Llama2-13B inference. Note that we omitted the concurrent data copy of PhOS. R stall stands for restore stall time.
@@
The runtime overhead of tracing GPU read and write set.    Figure 18(a) and (b) report the runtime overhead of the runtime validator on different applications. We observe a relatively small slowdown of 1–12% for various workloads. The observable overhead is small because: (1) the additional check only happens when the kernel accesses the global memory, which is less frequent for high computational efficiency; and (2) the instrumented kernels are only a small portion of the overall kernels in the workloads. Figure 18(c) reports the number of instrumented kernels invoked during the concurrent checkpoint and restore: We see that most of the kernels are not instrumented, e.g., in Llama2-13B inference, only 12% of the invoked kernels are instrumented.
Figure 18. The analysis of runtime validator overhead on (a) training and (b) inference workloads, along with (c) the ratios of kernels instrumented with the validator. Training is abbreviated as -T and inference as -I.
@@
Figure 19. The analysis of the copy-on-write overhead with different on-device buffer sizes.
Figure 19 shows the impact of the on-device buffer size on the checkpoint overhead when checkpointing a Llama2-13B training workload. We choose this workload because it has the largest memory footprint so little room is left for the on-device buffer (see Table 4). We report the increased stall time caused by the checkpoint, where the time is normalized to the stall time with our maximum configurable on-device buffer size (2 GB). We see that checkpoint overhead generally decreases with the on-device buffer size, but the difference is small, e.g., even without on-device buffer, PhOS incurs up to 9.1 % more stalls. This is because not all copy-on-write operations require stalling the kernel, so the slow copy through PCIe with insufficient on-device buffer can be hidden.
@@
Figure 20. Impact of checkpoint timing on the performance of Llama2-13B training. The blank in App is the stall caused by the CoW. The red indicates the data copy time. Note that we omitted the concurrent data copy of PhOS.
Figure 20 analyzes how the checkpoint timing affects the performance of PhOS on the CoW protocol. We omit analysis of the recopy protocol as it is similar. We control the timing with our SDK (see §A.2). We run Llama2-13B training workload and choose two times: (1) at the beginning of an iteration (before the forward pass), (2) at the end of the iteration (at the update phase). We measure the performance specifically on the 32th iteration to fully warm up the application. We can see that checkpoint at (1) is more efficient because few buffers are updated at the beginning: only activation buffers are updated (see also §8.5). Thus, if we can finish the checkpoint before reaching the update pass, we meet a few stalls due to CoW. In (1), copying 2.3 GB of data via 32 GBps PCIe [^1] The measured bandwidth is 25 GBps, slightly below the hardware limit. takes 185 ms, while the iteration time is 6.3 s.
@@
Figure 21. The heatmap of the read and write sets traced by our validation method. Workload: Llama2-13B training.
Figure 21 reports the GPU accesses traced by PhOS on Llama2-13B training. First, we can see that though PhOS traces accesses at the buffer level, it is still fine-grained enough to support an efficient CoW and recopy, because GPU applications typically allocate and write buffers in a fine-grained way. Moreover, the access patterns differ across different phases of execution, so the checkpoint timing is important.
Table 3. The success rate of our speculation-based GPU read and write sets tracing beyond the evaluated applications.
GPU Apps #Kernels / #Failures #Instances / #Failures
@@ Acknowledgment
We sincerely thank our shepherd Ashvin Goel and the reviewers from SOSP’24, OSDI’24 and SOSP’25 for their insightful feedback. We are grateful to Zhiyuan Dong, Dingyan Zhang and Xiating Xie for their valuable comments on the early draft of this paper. We also thank Shaoxun Zeng for sharing details on adapting PhOS to support CUDA graph. This work was supported in part by the National Key Research & Development Program of China (No. 2022YFB4500700), the National Natural Science Foundation of China (No. 62202291 and 62272291), as well as a research grant from Huawei.
@@ Appendix A Appendix
ResNET-152M PPO-336M Stable-Diffusion-1B Llama2-13B Llama3.3-70B
[…]
Application Training Inference Training Training Inference Training Inference Inference
@@
#GPUs 1 1 1 8 1 8 1 8
Memory usage (per GPU) 1.8 GB 1.7 GB 5.9 GB 70.6 GB 8.9 GB 73.6 GB 55.4 GB 70.8 GB
#Buffers (per GPU) 209 195 75 445 234 413 347 718
#Kernels (active) 13 8 41 51 50 36 74 73
@@
Problem formulation.    Consider a distributed computing job using $N$ GPUs, where each GPU fails $F$ times per hour. Assume these GPU failures are independent and identically distributed, following a uniform distribution over the entire computation interval ($T$). with PhOS, the checkpoint overhead is $O$. When a GPU fails, PhOS stops all GPU processes and restores from the latest checkpoint, with a restore time of $R$ for all GPUs.
@@
The total GPU time wasted due to failures at a checkpoint frequency of $f$ times per hour is:
$\Bigl{(}N\,F\,T\Bigr{)}\;\times\;\Bigl{(}R\;+\;\tfrac{N}{2f}\Bigr{)}\,.$
Put it all together.    The total GPU hours wasted due to checkpoint overhead and failure recovery is:
$\Bigl{(}N\,F\,T\Bigr{)}\;\times\;\Bigl{(}R\;+\;\tfrac{N}{2f}\Bigr{)}\;+\;N\,O\,f\,T\,.$
Solving the optimal $f$.    For a given job, the frequency of failures ($F$), the checkpoint overhead ($O$), and the restore time ($R$) are static and can be profiled online (DBLP:conf/fast/MohanPC21, ). Therefore, we only need to determine the optimal checkpoint frequency $f$ to minimize wasted GPU hours. By differentiating the total GPU hours wasted with respect to $f$, we derive:
@@
Figure 22. An illustration of how applications can use the PhOS SDK to control checkpoint timing. Lines marked with “+” indicate code that interacts with PhOS.
Applications can explicitly call our SDK to control checkpoint timing and the protocol used, as shown in Figure 22. Our OS-level design merely requires minimal code changes—just one line of code. For example, the code instructs PhOS to checkpoint the process at the beginning of a training iteration. Since the checkpoint is executed asynchronously, it will not block application execution unless the last checkpoint is still in progress.
@@
Table 4 shows the detailed buffer allocation information as well as the number of GPU kernels information of our evaluated applications. For training workloads: ResNET uses the CIFAR-10 dataset with a batch size of 32; PPO uses the gym code (gym, ); Stable Diffusion uses the SD v1-4 model with a batch size of 1,536 per GPU; and Llama uses a distributed configuration of 8TP (Tensor Parallelism) and 1DP (Data Parallelism) with a batch size of 4 due to GPU resource constraints. All training workloads use the AdamW optimizer. We use the same training datasets to evaluate the inference workloads.